### 24 Jan 2024 v2

# bibliography

[…]
Supplementary information
Prompts
Intuitive physics
For the intuitive physics experiment, we used a task from Lerer et al. [lerer2016learning]. For each trial, we asked the models three questions:
Question 1
What is the background color?
Question 2
How many blocks are in the image?
Question 3
Will this block tower fall? Give a boolean answer.
Causal reasoning
For the causality experiment, we used a task from Zhou et al. [zhou2023mental, gerstenberg2017faulty]. For each trial, we again asked the models three questions:
Question 1
How many blocks are there in the image? You are only allowed to answer with a single number. No words allowed!
Question 2
How many of the red bricks would fall off the table if the dark grey brick wasn’t there? You are only allowed to answer with a single number between 0 and {num_blocks} corresponding to how many blocks would fall. No words allowed!
Question 3
How responsible is the dark grey brick for the red bricks staying on the table? You are only allowed to answer with a number on a scale from 0% (not at all responsible) to 100% (fully responsible). No words allowed!
Naive Utility Calculus
For the intuitive psychology experiment, we used three tasks from Jara-Ettinger et al. [jara2020naive]. We again asked the model two descriptive questions in order to assess their basic comprehension of the scene:
Question 1
Please answer the following question with a single color name only: What color is the background of the central part of the image? You are only allowed to answer with a single color name!
Question 2
Please answer the following question with a number only: How many orange or white containers are in the image? You are only allowed to answer with a number!
For experiment 1A we first gave the models the following basic prompt, which was combined with different trial specific questions:
[…]
This task is about astronauts. The astronauts are exploring planets with alien terrains depicted with different colours and textures. Each astronaut has different skills, making each terrain more or less exhausting or easy for them to cross. All astronauts can ultimately cross all terrains, even if it’s exhausting. The astronauts land far from the base and have to walk there. In each image, the black circle on the left indicates where the astronaut landed. The base is on the middle right part of the image. Sometimes care packages are dropped from above and the astronauts can pick them up. There are two kinds of care packages depicted with an orange cylinder and a white cube. Each astronaut has different preferences and likes each kind of care package in different amounts. The astronauts don’t actually need the care packages. They can go straight to the base, or they can pick one up. You will see images of different astronauts with different skills and preferences travelling from their landing location to the home base. The astronauts always have a map. So they know all about the terrains and the care packages. Please answer the following question with a number only:
For experiment 1B we first gave the models the following basic prompt, which was again combined with different trial-specific questions:
[…]
This task is about astronauts. The astronauts are exploring planets with alien terrains depicted with different colours and textures. Each astronaut has different skills, making each terrain more or less exhausting or easy for them to cross. All astronauts can ultimately cross all terrains, even if it’s exhausting. Sometimes, the astronauts land far from the base and have to walk there. In each image, the black circle indicates where the astronaut landed. The base is in the center of the image. Sometimes care packages are dropped from above and the astronauts can pick them up. There are two kinds of care packages depicted with an orange cylinder and a white cube. Sometimes both care packages are identical. The astronauts cannot pick both care packages. Each astronaut has different preferences and likes each kind of care package in different amounts. The astronauts don’t actually need the care packages. They can go straight to the base, or they can pick one up. You will see images of different astronauts with different skills and preferences travelling from their landing location to the home base. The astronauts always have a map. So they know all about the terrains and the care packages. Please answer the following question with a number only:
Finally, for experiment 1C we first gave the models the following basic prompt:
[…]
This task is about astronauts. The astronauts are exploring planets with alien terrains depicted with different colours and textures. Each astronaut has different skills, making each terrain more or less exhausting or easy for them to cross. All astronauts can ultimately cross all terrains, even if it’s exhausting. The astronauts land far from the base and have to walk there. In each image, the black circle on the left indicates where the astronaut landed. The base is on the right part of the image. The path astronauts take from where they land to their base is indicated by a thick black line between the black circle on the left and the astronaut on the right. Sometimes care packages depicted by a blue cube on a black background are dropped from above and the astronauts can pick them up. Each astronaut has different preferences and likes each care package in different amounts. The astronauts don’t actually need the care packages. They can go straight to the base, or they can pick one up. You will see images of different astronauts with different skills and preferences travelling from their landing location to the home base. Your task is to judge how easy/exhausting it is for the astronaut in each image to cross each terrain, and how much they like each care package. The astronauts always have a map. So they know all about the terrains and the care packages. Please answer the following question with a number only:
Experiment 1A
Images Questions
0, 1
  1. 1.
    How easy is it for the astronaut in this image to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How much does the astronaut in this image like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
2, 3, 4, 5
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!"
6, 7, 8, 9, 10
  1. 1.
    How easy is it for the astronaut to cross the purple terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
11, 12, 13, 14, 15
  1. 1.
    How easy is it for the astronaut to cross the purple terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  4. 4.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!"
Experiment 1B
Images Questions
1, 2, 3, 4, 5, 6, 7
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the yellow terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  4. 4.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
8, 9, 10
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the yellow terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
11, 12, 13, 14, 15, 16, 17
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
Experiment 1C
Images Questions
All images
  1. 1.
    How easy is it for the astronaut to cross the yellow terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the blue care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!